A goldsmith mixes 10oz of a 30% gold alloy with 20oz of a 15% gold alloy. what is the percent concentration of the resulting alloy?
Answer by richwmiller(17219)

10*.30+20*.15=30*x
3+3=30x
6=30x
divide by 6
1=5x
.20=x
20%=x